# Capturing Human Motion from Monocular Images in World Space with Weak-supervised Calibration

### v4

@@ V Discussion
Future Another vital direction for human motion recovery in world space is to capture the global trajectory of 3D human bodies. W-HMR mainly solves the problems caused by camera position and rotation. Although W-HMR can recover the human body posture in world space, it cannot obtain the absolute spatial position of the human body. Absolute spatial position is essential for the practical application of motion capture. At present, most related research still needs to rely on additional equipment to assist. Some methods that can predict 3D pose [63, 51, 64] or SMPL [65, 66] in the world coordinate from monocular images have been proposed but still far from practical. This will be an important future research direction for motion capture based on monocular images.
VI Conclusion
We proposed the first model capable of reconstructing human bodies accurately and reasonably in both camera and world coordinates. To achieve this, we designed a weak-supervised camera calibration method. ”Weak-Supervised” solves the problem of insufficient focal length labels. ”Camera calibration” solves the problem caused by image distortion. The focal length predictor can generate appropriate focal lengths based on body distortion information, eliminating the dependence on background information and focal length labels, which are more suitable for human recovery tasks.

### v6

@@ 5 Discussion
Future Another vital direction for human motion recovery in world space is to capture the global trajectory of 3D human bodies. W-HMR mainly solves the problems caused by camera position and rotation. Although W-HMR can recover the human body posture in world space, it cannot obtain the absolute spatial position of the human body. Absolute spatial position is essential for the practical application of motion capture. At present, most related research still needs to rely on additional equipment to assist. Some methods that can predict 3D pose (Moon et al., 2019; Mehta et al., 2017b; Shimada et al., 2020) or SMPL (Sun et al., 2022; Tripathi et al., 2023) in the world coordinate from monocular images have been proposed but still far from practical. This will be an important future research direction for motion capture based on monocular images.
6 Conclusion
We proposed the first model capable of reconstructing human bodies accurately and reasonably in both camera and world coordinates. To achieve this, we designed a weak-supervised camera calibration method. “Weak-Supervised” solves the problem of insufficient focal length labels. “Camera calibration” solves the problem caused by image distortion. The focal length predictor can generate appropriate focal lengths based on body distortion information, eliminating the dependence on background information and focal length labels, which are more suitable for human recovery tasks.